OpenStudy (anonymous):

A building has a ramp to its front doors to accommodate the handicapped. If the distance from the building to the end of the ramp is 17 feet and the height from the ground to the front doors is 7 feet, how long is the ramp? (Round to the nearest tenth.)

OpenStudy (dumbcow):

use pythagorean thm a^2 +b^2 = c^2

OpenStudy (anonymous):

7^2+b^2=17^2?